OpenStudy (anonymous):

This is probably something dumb, But What is wrong here? x=0, Differenciating w.r.t. x, => 1 = 0

hartnn (hartnn):

u differentiate a function, not an equation.

hartnn (hartnn):

f(x)=x f'(x)=1 f()=0 f'()=0

OpenStudy (anonymous):

x=0 is the y axis. It's basically a straight line.

hartnn (hartnn):

or in other words.. d/dx(x) = d/dx (0) dx/dx=0 dx=0

hartnn (hartnn):

right, when u differentiate at a point, u find slope of that line.

OpenStudy (anonymous):

precisely. Show should I be getting infinity?

OpenStudy (anonymous):

*So shouldn't

hartnn (hartnn):

u cannot write vertical line as f(x)=.....

OpenStudy (anonymous):

I see. That kinda makes sense. Thankyou :)

hartnn (hartnn):

really! ? but u got me thinking.....

OpenStudy (anonymous):

Yeah. I think I get it. Because you used the word function. x=0 is not a function of x at all. Therefore it cant be differenciated w.r.t. x. Reason being for the same value of x (=0), you have multiple values of y. Therefore not a function => Limit of the differential doesnt exist (Using calculus definition) Also, to support this, We can say that x=0 is a function of y. As for one value of y you get one x. Therefore differential w.r.t y should exist. And note here, dx/dy = 0which is what should have been.

OpenStudy (anonymous):

@hartnn Does that convince you?

hartnn (hartnn):

oh! yesss.....that 'x=0 is not a function' part! why didn't it click me?? i knew that....

hartnn (hartnn):

thanks @siddhantsharan :D